the
company
Old
Co has a
current stock
price of $210 per share
. L
ast year
they
paid an
annual dividend of $10 per share. Historically the dividend growth was 5% per year
;
the next
dividend
payment will be $
1
0.50
. What is the discount rate implied by this firm?
1
Expert's answer
2020-04-24T11:14:21-0400

The formula for share price is:"P = D0*(1 + g)\/(r - g)",

"r = D0*(1 + g)\/P + g = 10*1.05\/210 + 0.05 = 0.1."